BLAST SHIFTS RESUME TOMORROW (WEDNESDAY) MORNING AT 9 AM!!!!!
Wire Chamber gas:  The return lines have been split, and the L and R flows 
roughly balanced by watching the bubblers.  The valve in the left supply 
line is open slightly more than we had it on the weekend.  If the valve is 
open all the way, the left bubbles FAR more than the R (10 times?).  Jim 
and Chris found a leak in the flexible line for the LEFT supply.  There is 
also a crimp in the right return.  They will be making up or buying new 
flexible lines, which might be ready early next week.  The earliest higher 
flow rate MFCs can be shipped is Friday.  Until then we will run 5 l/m of 
helium, and hope that the balanced flow helps both L and R drift time 
spectra.  The helium bottles have been changed, but the isobutane has not.
ABS:  The dissociator tube has been replaced.  Genya will be in this 
evening to work on tuning the hydrogen transitions with the BLAST and 
holding fields on.
Beam:  All areas except the South Hall will be searched overnight, and beam 
will be run to the 14 degree area.  Tomorrow morning we'll remove the wire 
chamber covers, then search the SH.  Then we'll run beam on unpolarized 
hydrogen and test the chambers.
Solenoids: Filled today.
                                        Karen
This archive was generated by hypermail 2.1.2 : Mon Feb 24 2014 - 14:07:31 EST